OpenStudy (anonymous):

The typical American diet has too much salt (NaCl) in it. A healthy adult should limit his or her salt intake to 1 tsp of salt maximum per day, which equals about 6 g. b. How many moles should an adult's maximum intake be? (2 points) c. How many molecules of NaCl is the daily maximum? (2 points)

OpenStudy (aaronq):

for a) use: \(\sf moles=\dfrac{mass ~(in~g)}{Molar ~mass~(in~g/mol)}\) for b) use: \(\sf moles=\dfrac{particles }{Avogadro's~number}\) particles in this case are molecules

OpenStudy (anonymous):

ok so can you check my work?

OpenStudy (anonymous):

What is the molar mass of NaCl? (2 points) b. How many moles should an adult's maximum intake be? c. How many molecules of NaCl is the daily maximum? d. What are the percent compositions of Na and Cl in NaCl? (Round your answers to the nearest tenth percentage.) e. Processed foods often contain large amounts of salt, but it is actually sodium (Na) that causes health concerns. If 6 g of salt (NaCl) should be an adult's daily maximum, what would the daily maximum of just sodium (Na) be? Give your answer in milligrams. f. The sodium (Na) content of 1 particular can of soup is 890 mg sodium in 1 c. What is the molarity of sodium (Na) in the soup? (Note: 4 c = 1 qt, and 1 qt = 0.946 L.)

OpenStudy (aaronq):

post one question at a time please.

OpenStudy (anonymous):

so for a. 58.55 g/mol d. Na = 39.3 percent, Cl = 60.7 percent

OpenStudy (anonymous):

sorry okay so first question

OpenStudy (anonymous):

What is the molar mass of NaCl? I got 58.4 g/mol

OpenStudy (anonymous):

can you check thanks!

OpenStudy (aaronq):

yep thats the molar mass, the question asks for moles though

OpenStudy (aaronq):

oh wait sorry, thats b).

OpenStudy (anonymous):

nope the question asks for molar mass okay :)

OpenStudy (anonymous):

i dont understand b and c though

OpenStudy (anonymous):

so would i do 6g divided

OpenStudy (anonymous):

*divided by 58.55 g/mol

OpenStudy (aaronq):

yep thats it

OpenStudy (anonymous):

what about c though..

OpenStudy (aaronq):

use the second equation i posted using the moles you get from part b) ps. Avogadro's number is a constant

OpenStudy (anonymous):

i know avogrado's number = 6.02

OpenStudy (anonymous):

well 6.02 x 10^23

OpenStudy (anonymous):

but what do you mean by particles?

OpenStudy (anonymous):

0.103*6.023*10^23

OpenStudy (anonymous):

got it !

OpenStudy (aaronq):

thats it

OpenStudy (anonymous):

okay so i got answers for e, f, and g

OpenStudy (anonymous):

can you check them?

OpenStudy (anonymous):

e= 2359 mg, f= 0.2 m NA

OpenStudy (anonymous):

sorry just f and g

OpenStudy (aaronq):

there is no g)

OpenStudy (anonymous):

yea i know i meant just e and f sorry !

OpenStudy (anonymous):

sorry i have this huge chem assignment due tomorrow and i kinda fell asleep in class ahh

OpenStudy (aaronq):

both are incorrect

OpenStudy (anonymous):

wait can you correct me then? for part e, i did 39.3 percent times 6 g

OpenStudy (anonymous):

and for part f i did 890/22.989 g/mol = 0.0387 moles i converted to 0.2365 L and i did molarity = 0.0387 / 0.2365 L = 0.2

OpenStudy (aaronq):

Oh wait e) is fine. For f, you forgot to convert mg to g.

OpenStudy (anonymous):

oh and here are the questions again: The typical American diet has too much salt (NaCl) in it. A healthy adult should limit his or her salt intake to 1 tsp of salt maximum per day, which equals about 6 g. Use this information to answer the following questions. (Round atomic masses to the nearest tenth unit.) E. Processed foods often contain large amounts of salt, but it is actually sodium (Na) that causes health concerns. If 6 g of salt (NaCl) should be an adult's daily maximum, what would the daily maximum of just sodium (Na) be? Give your answer in milligrams. F. The sodium (Na) content of 1 particular can of soup is 890 mg sodium in 1 c. What is the molarity of sodium (Na) in the soup? (Note: 4 c = 1 qt, and 1 qt = 0.946 L.)

OpenStudy (anonymous):

okay so e is 2358 mg

OpenStudy (anonymous):

wait no i did f right

OpenStudy (anonymous):

i'll post a file hold on.

OpenStudy (anonymous):

i did do g

OpenStudy (aaronq):

nope, your mass should be in grams. Otherwise the units in the molar mass wouldn't cancel.

OpenStudy (anonymous):

wait a sec. i have a copy written

OpenStudy (aaronq):

you converted mg to grams in this, but not in what you posted. so it's right

OpenStudy (anonymous):

so is 0.2 m Na right?

OpenStudy (aaronq):

yes, use a capital M for molarity as lowercase m means molality

OpenStudy (anonymous):

oh okay. thank you so much. most helpful responses!

OpenStudy (anonymous):

if you don't mind, i have a few more questions

OpenStudy (aaronq):

no problem! post 1 question per thread though

OpenStudy (anonymous):

okay